# Lee-Wick instability; pion condensate; strangelets; strange stars

### v1

@@ Abstract
Researchers have long been interested in exotic states of matter. In the early 1970s, Migdal proposed existence of metastable or stable nuclei containing a pion condensate, and Bodmer posited existence of collapsed nuclei with a quark core. Lee and Wick proposed that compressed nuclear matter could undergo a phase transition into a scalar condensate state with effectively very light baryons. In the mid-1980s, Witten pointed out that strange quark matter may be an absolutely stable form of matter that can exist in the form of quark strangelets. These hypotheses stimulated large experimental efforts to detect these states in heavy-ion collisions and in terrestrial samples. Additionally, the cosmos is the arena for a broad search for traces of exotic matter in compact objects and other energetic phenomena. We review various approaches and arguments for a possibility of abnormal states of matter, including metastable and stable dense objects with condensates, as well as dilute nuclear objects. Among these, we discuss p-wave and s-wave pion condensations in nuclear systems and formation of a $\Delta$ resonance matter. Then we discuss a scalar mode condensation, which is triggered by the Pomeranchuk instability in dilute nuclear matter, and a clustering. We also mention recent ideas that a relativistic rotation and a dark-matter component may stabilize abnormal nuclear systems. Finally, we list some observational anomalies which cannot be yet appropriately explained by conventional physics.
keywords:
Lee-Wick instability; pion condensate; strangelets; strange stars
@@
1 Introduction
A search for abnormal forms of matter accompanied the development of modern physics during the 20th century. In year 1931 Lev Landau wrote a paper Landau (1932), in which he derived an upper limit of star mass consisting of a ultra-relativistic gas $\sim 1.5\,M_{\odot}$ and hypothesized that heavier stars may possess regions in which the laws of quantum mechanics (and therefore of quantum statistics) are violated. In such “unheimliche Sterne” (Germ eerie stars) as they were dubbed by Landau, see Yakovlev et al. (2013), electroneutral mixture of protons and electrons, the only particles known by then, would be stable against particle annihilation. The discovery of neutron by James Chadwick in 1932 lead Dmitri Ivanenko to the idea Iwanenko (1932) that nuclei may consist of protons and neutrons instead of protons and electrons as thought before. In Heisenberg (1932a, b, 1933) Werner Heisenberg introduced the theory of an atomic nucleus consisting of nucleons, — strongly interacting particle with the putative isospin 1/2, which manifest themselves as protons and neutrons depending of the nucleon isospin projections. In 1934 Walter Baade and Fritz Zwicky Baade and Zwicky (1934) proposed that sufficiently massive stars end their life undergoing a transition (supernova event) into compact object consisting mainly of neutrons and called, therefore, neutron stars. More than three decades the neutron stars were considered as a theoretical exotica. Only 1967 Jocelyn Bell and Anthony Hewish reported Hewish et al. (1968) the first observation of a radio pulsar. By now neutron stars are not considered as exotica, and more than 3000 pulsars are observed, presently.
What is the ground state of the world around us has been one of the ‘big’ questions since the beginning of physics and then at the discussion of the role of the either. If there is some underlying symmetry behind the construction of our world, would it be possible to change the ground state? In condensed matter physics this concept was proposed by Landau in 1937 Landau (1937a, b) for the description of phase transitions. He pointed out that any phase transition may be characterized by a change in the value and/or the symmetry structure of some order parameter, which can be used as an effective degree of freedom in the thermodynamic potential. This order parameter can have a rather complicated relation to the fundamental degrees of freedom of the considered system. Generalization to the case of an order parameter interacting with the magnetic field was done by Vitaly Ginzburg and Lev Landau in 1950 Ginzburg and Landau (1950).
Addressing the question of a possible transformation of the ground state of our physical world it is natural to suggest a simplest choice that the order parameter should have quantum number of vacuum $0^{+}$ (spin and parity), i.e., correspond to some scalar field. There were no scalar particles known in the mid 60s. Nevertheless, Julian Schwinger tried to formulate a theory of strong interactions using the spin 0, spin 1, and spin 1/2 fields Schwinger (1957). The first attempts to build a model of interacting hadrons using a symmetry between vector and axial currents, which enter equally in weak interactions, were undertaken in works by John Polkinghorn, Murrey Gell-Mann and Maurice Lévy Adler and Bardeen (1969); Gell-Mann and Levy (1960) where the so-called $\sigma$-model, including scalar and pseudoscalar mesons coupled to nucleons, was formulated. The Lagrangian of this model contains the part symmetric with respect to local transformations keeping parity of the states and transformations mixing states with different parities, (the chiral invariant part) and the part, which is not invariant under the parity mixing transformations. Our physical ground state ($0^{+}$) is chosen as a result of the spontaneous symmetry breaking in the invariant part of the Lagrangian whereas the symmetry breaking part is treated as a rather weak perturbation. A dynamical mechanism of the nucleon mass generation through the mechanism of spontaneous breaking of the chiral symmetry was elaborated in independent works by Valentin Vaks and Anatoly Larkin Vaks and Larkin (1961) and Yoichiro Nambu Nambu (1960), applying superconductivity methods to a Heisenberg type four-fermion two-component Lagrangian.
We should certainly mention a role played by the spontaneous symmetry breaking in the unified electroweak theory of Abdul Salam, Sidney Glashow, and Steven Weinberg and the role played by the Higgs scalar boson in it Glashow (1959, 1961); Salam and Ward (1959, 1961); Weinberg (1967). The study of the properties of superdense matter described by unified gauge theories began in 1972 with the work of David Kirzhnits Kirzhnits (1972), who showed that the classical scalar field responsible for symmetry breaking disappears at a high enough temperature. In 1974 first grand unified theory was proposed by Howard Georgi and Sidney Glashow Georgi and Glashow (1974).
Now let us come to the subject of this review, discussion of yet considered, as exotica, possibilities of existence of various abnormal nuclear systems. Soon after prediction of quarks by Murrey Gell-Mann and George Zweig in 1964, Dmitry Ivanenko and Dmitry Kurdgelaidze suggested that a neutron star may transform in a quark star Ivanenko and Kurdgelaidze (1965), see also Itoh (1970). In 1971 Arnold Bodmer proposed that also atomic nuclei may undergo a transition in a superdense quark state Bodmer (1971).
In 1971 Arkady Migdal put forward the idea that ground state of atomic nuclei may contain a pseudoscalar condensate ($0^{-}$) of pions. He demonstrated occurrence of the pion condensation in a sufficiently deep potential well in scalar and electric fields forming deep potential wells for pions and suggested possibility of superdense stable or metastable nuclei glued by the pion condensate Migdal (1971). Three types of abnormal $\pi$ condensate nuclei: superdense, neutron and supercharged ones were discussed in Migdal (1974); Migdal et al. (1976), more details see in reviews Migdal (1978); Migdal et al. (1990).
In 1974 Tsung-Dao Lee and Gian-Carlo Wick suggested possible existence of a superdense state of nuclear matter characterized by a value of the scalar (e.g., $\sigma$) field different from its mean value in atomic nuclei Lee (1974); Lee and Wick (1974). This idea was reviewed in Lee (1975, 1976). The simplest variant of the model contains only two fields, the neutral scalar meson field and the isospinor nucleon field. As the $\pi$ condensate nuclei, the $\sigma$ abnormal nuclear state can be either stable or metastable. In the former case the normal nuclei after a while would undergo the phase transition to the new state. In first works the mean field approximation was used in spite of the assumed strong interaction of the scalar meson with the nucleons. To somehow justify this assumption the authors noted analogy between the abnormal state and the Bose–Einstein condensation one. Then Lee and Margulies argued Lee and Margulies (1975); Lee and Wick (1979) that quantum fluctuations may not change the statement about the existence of a super-dense state of nuclear matter.
At the workshop at Bear Mountain, to north from New York City, on “BeV/nucleon collisions of heavy ions” Lee presented his results on the possible vacuum reconfiguration Lee (1974) and noted Lederman and Weneser (1975) that “in order to produce the abnormal nuclear state, we must consider reactions in which (i) a large number of nucleons is involved and (ii) the nucleon density can be increased by a significant factor. One is, therefore, led to considerations of high energy collisions between heavy ions, say ${\rm U}+{\rm U}\to Ab+\dots$, where $Ab$ denotes the abnormal nuclear state.” The event at Bear Mountain was pivotal in the conception of heavy ion physics. 1974 there were the Synchrophasotron at the Joint Institute for Nuclear Research in Dubna and a high-energy heavy-ion accelerator, the Bevalac, being completed at the Lawrence Berkeley Laboratory in Berkeley. Ideas about possibility of the production of abnormal nuclei in heavy-ion collisions gave a push to active experimental investigations. One suggested also the look for collapsed quark nuclei Bodmer (1971) and superdense nuclei with pion condensate in heavy-ion collisions, see Refs. Aleshin et al. (1976); Galitskii (1976). A possibility of reaching high degree of compression of nuclear matter in shock waves that could help to from anomalously dense droplets in heavy-in collisions was investigated in Refs. Scheid et al. (1974); Sobel et al. (1975). Limits on the occurrence of abnormal nuclei in heavy-ion collisions and on the terrestrial concentration of abnormal nuclei were discussed Holt et al. (1976); Karnaukhov and Polikarpov (1977).
Cosmos was another domain to look for an anomalous state of matter. Yakov Zeldovich liked to claim that the Universe is a poor man’s accelerator: experiments don’t need to be funded, and all we have to do is to collect the experimental data and interpret them properly. Years later ideas of spontaneous symmetry breaking penetrated in cosmological models, cf. Linde (1990).
Probably in 1975, Migdal returned from a conference and told one of us (D.N.V.) that during his talk, in which he mentioned a possibility of the existence of anomalous supercharged nuclei with a pion condensate, Gell-Mann asked him about a role of electrons in such highly-charged systems. Migdal suggested to think about this problem. In works Migdal et al. (1976, 1977) the authors developed a relativistic semiclassical treatment of the Dirac equation in deep electric potentials and the many-particle approximation for electrons occupying levels of the lower continuum. The occupation of the levels by electrons and negative muons in the deep electric potential well was named “the fermion condensation”. The presence of the upper bound on the electric potential was discussed by Berndt Müller and Johann Rafelski in Müller and Rafelski (1975). Preparing this manuscript, we realized that Lee also discussed a role of electrons in abnormal nuclei. He wrote in Ref. Lee (1976) that “…the abnormal state can create $e^{+}e^{-}$ pairs. The $e^{+}$ will be sent to infinity, but a fair fraction of the $e^{-}$ will be kept within the abnormal nucleus. As $Z$ increases, the number of $e^{-}$ also increases. The interplay between the added Fermi energy of $e^{-}$ and the Coulomb energy may eventually bring the abnormal state to the point of instability when $A$ reaches $\approx 10^{4}$.” References Voskresensky et al. (1977); Voskresensky and Chernoutsan (1978); Voskresensky (1977) described the charged condensate pions and vacuum electrons in strong electric fields and suggested a possibility of existence of nuclei-stars of arbitrary size (up to the compact star size) glued by the charged $\pi$ condensate produced in the dense nuclear matter and the electrons produced from the vacuum.
In 1979 Mikhail Troitsky and Nikolai Chekunaev discussed in Troitsky and Chekunaev (1979) that at densities above the nuclear saturation density $n_{0}=0.16$ fm-3 the gain in the energy due to the pion condensation can exceed the energy loss associated with the transformation of nucleons into heavier baryon species like hyperons and $\Delta$ isobars. The similar idea was applied to quarks by Edward Witten in 1984. He conjectured in Witten (1984) that at increasing baryon density it becomes energetically favorable to occupy the $s$ quark Fermi sea. He argued that “strange matter” containing comparable numbers of up, down, and strange quarks may become absolutely stable. Edward Farhi and Robert Jaffe studied the strange matter with intermediate baryon number $10^{2}<A<10^{7}$ Farhi and Jaffe (1984), later Alvaro De Rujula and Sidney Glashow De Rujula and Glashow (1984); De Rujula (1985) suggested a possibility of existence of quark nuggets, larger size nuclearites and strange stars. The charge distribution in the quark nuggets and strange stars was considered in Alcock et al. (1986) in a similar manner as was done for supercharged nuclei in Refs. Migdal et al. (1977); Voskresensky et al. (1977); Voskresensky and Chernoutsan (1978). It was suggested that there may exist stars with hadronic shell and quark core –hybrid stars Shapiro and Teukolsky (1983); Bethe et al. (1987). First investigations were done ignoring possibility of color superconductivity. As a detailed review one may use Madsen (1999). Presence of the Cooper pairing in quark matter was known starting from works Barrois (1977); Bailin and Love (1984), where gaps were estimated as $\Delta\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}$ MeV, as for the nucleons. Then interest to possibility of quark stars was renewed after arguments were given in Rapp et al. (1998); Alford et al. (1998) in favor of large values of the quark-quark pairing gaps, $\Delta\sim 100$ MeV, see Alford et al. (2008) and references therein.
The axion quark nuggets (AQN) model was proposed in Refs. Zhitnitsky (2003, 2021); Zhitnitsky and Maroudas (2025). The latter model incorporates axion domain walls formed during the QCD phase transition. The domain walls act as stabilizing structures, alleviating the need for a first-order phase transition as in the original Witten model. A generic feature of the quark nuggets is the electrosphere. In case of antimatter axion quark nuggets the electrosphere of the $\sim 10^{-8}$ cm depth contains positrons. The axion quark nuggets behave as conventional dark matter components in the low density environment and become strongly interacting macroscopically large objects in the relatively high-density environment. The available energy due to matter-antimatter annihilation can be as high as $\sim 2$ GeV per baryon charge.
[…]
@@ 1 Introduction
Although the initial studies Lee (1974); Lee and Wick (1974); Migdal (1974) discussed the possibility of the sigma or pion condensation occurring at densities both greater than and less than $n_{0}$, it soon became evident that the latter possibility was unrealistic. On the other hand various models of equation of state of the dilute nearly isospin-symmetric nuclear matter demonstrate the presence of region of the first-order phase transition and spinodal instability at nucleon densities $0.3n_{0}\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}n\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}0.7n_{0}$ Röpke et al. (1982); Curtin et al. (1983); Bertsch and Siemens (1983); Schulz et al. (1983); Borderie et al. (2022). In this density region the compressibility of nuclear matter is negative and the scalar Landau-Migdal parameter of the Fermi liquid $f_{0}(n)$ is smaller than 1-1, that causes the Pomeranchuck instability. In Kolomeitsev and Voskresensky (2016) it was shown that the Pomeranchuck instability may lead to a condensation of an effective scalar field in a Fermi liquid. The presence of such condensate would remove instabilities in first and zeroth sounds. Owing to the condensate, a novel metastable state might exist at nuclear subsaturation density for $N\sim Z$, where $N$ is the number of neutrons and $Z$ is the number of protons. In Maslov and Voskresensky (2019) it was realized that the stability region remains also in case of significantly isospin-asymmetric matter.
@@
It was found in Refs, Voskresensky (2024, 2025) that a rapidly rotating drop of nuclear matter, being not bound in the rest frame, may become bound in the rotation frame because of the formation of a charged pion condensate giant vortex. This observation invites looking for the nuclear drops with high angular momenta.
The existence of stable or metastable nuggets would have numerous important consequences in physics. Therefore, it would be worthwhile to continue theoretical and, especially, experimental investigations of such objects, although all of the above predictions are based on speculation about the possible behavior of dense hadron and quark matter that has not been tested against experimental data. In the body of this review we focus on some more specific details.
The paper has the following structure. In Sec. 2 we focus on the Lee-Wick model. Section 3 considers various aspects of the model of supercharged nuclei glued by light bosons. In Sec. 4 we discuss superheavy charged pion condensate nuclei, then $\pi^{0}$ ferromagnetism, and then we pay attention to subtleties associated with description of the $s$ wave pion-nucleon interaction and $s$ wave pion condensation. Section 5 discusses the $\Delta$ resonance matter. In Sec. 6 we focus on the hypothesis of strangelets. Section 7 presents the idea of existence of the dilute metastable nuclear state glued by a scalar condensate field. Section 8 discusses anomalies that defy conventional physics interpretations, part of which could be related to abnormal nuclear systems. The last Sec. 9 contains concluding remarks.
@@
2 The Lee-Wick model
Let us present the idea of Lee and Wick in a nut shell. Let us consider the Lagrangian density of nucleons interacting with a scalar field $\phi$:

### v2

# arXiv:2509.19848v2[nucl-th] 02 Mar 2026

@@ Abstract
Researchers have long been interested in exotic states of matter. In the early 1970s, Migdal proposed existence of metastable or stable nuclei containing a pion condensate, and Bodmer posited existence of collapsed nuclei with a quark core. Lee and Wick proposed that compressed nuclear matter could undergo a phase transition into a scalar condensate state with effectively very light baryons. In the mid-1980s, Witten pointed out that strange quark matter may be an absolutely stable form of matter that can exist in the form of quark strangelets. These hypotheses stimulated large experimental efforts to detect these states in heavy-ion collisions and in terrestrial samples. Additionally, the cosmos is the arena for a broad search for traces of exotic matter in compact objects and other energetic phenomena. We review various approaches and arguments for a possibility of abnormal states of matter, including metastable and stable dense objects with condensates, as well as dilute nuclear objects. Among these, we discuss scalar condensate, p-wave and s-wave pion condensations in dense nuclear systems and formation of a $\Delta$ resonance matter. Then we consider a scalar mode condensation triggered by the Pomeranchuk instability in dilute nuclear matter, and a clustering. We also mention recent ideas that a relativistic rotation and a dark-matter component may stabilize abnormal nuclear systems. Finally, we list some observational anomalies which cannot be yet appropriately explained by conventional physics.
I Introduction
A search for abnormal forms of matter accompanied the development of modern physics during the 20th century. In year 1931 Lev Landau wrote a paper [173], in which he derived an upper limit of star mass consisting of an ultra-relativistic gas. He hypothesized that stars heavier than $\sim 1.5\,M_{\odot}$ may possess regions, in which the laws of quantum mechanics (and therefore of quantum statistics) are violated. In such “unheimliche Sterne” (Germ: eerie stars) as they were dubbed by Landau, see [300], electroneutral mixture of protons and electrons, the only particles known by then, would be stable against particle annihilation. The discovery of neutron by James Chadwick in 1932 lead Dmitri Ivanenko to the idea [133] that nuclei may consist of protons and neutrons instead of protons and electrons as thought before. In [119, 120, 121] Werner Heisenberg introduced the theory of an atomic nucleus consisting of nucleons, — strongly interacting particles with the putative isospin 1/2, which manifest themselves as protons and neutrons depending of the nucleon isospin projections. In 1934 Walter Baade and Fritz Zwicky [17] proposed that sufficiently massive stars end their life undergoing a transition (supernova event) into compact object consisting mainly of neutrons and called, therefore, neutron stars. More than three decades the neutron stars were considered as a theoretical exotica. Only in 1967 Jocelyn Bell and Anthony Hewish reported [122] the first observation of a radio pulsar. By now neutron stars are not considered as exotica, and more than 3000 pulsars are observed, presently.
What is the ground state of the world around us has been one of the ‘big’ questions since the beginning of physics and then at the discussion of the role of the either. If there is some underlying symmetry behind the construction of our world, would it be possible to change the ground state? In condensed matter physics this concept was proposed by Landau in 1937 [171, 172] for the description of phase transitions. He pointed out that any phase transition may be characterized by a change in the value and/or the symmetry structure of some order parameter, which can be used as an effective degree of freedom in the thermodynamic potential. This order parameter can have a rather complicated relation to the fundamental degrees of freedom of the considered system. Generalization to the case of an order parameter interacting with the magnetic field was done by Vitaly Ginzburg and Lev Landau in 1950 [100].
Addressing the question of a possible transformation of the ground state of our physical world it is natural to suggest a simplest choice that the order parameter should have quantum number of vacuum $0^{+}$ (spin and parity), i.e., correspond to some scalar field. There were no scalar particles known in the mid 60s. Nevertheless, Julian Schwinger tried to formulate a theory of strong interactions using the spin 0, spin 1, and spin 1/2 fields [257]. The first attempts to build a model of interacting hadrons using a symmetry between vector and axial currents, which enter equally in weak interactions, were undertaken in works by John Polkinghorn, Murrey Gell-Mann and Maurice Lévy [3, 96]. The so-called $\sigma$-model, including scalar and pseudoscalar mesons coupled to nucleons, was formulated. The Lagrangian of this model contains the part symmetric with respect to local transformations keeping parity of the states (the chiral invariant part) and the part, which is not invariant under the parity mixing transformations. Our physical ground state ($0^{+}$) is chosen as a result of the spontaneous symmetry breaking in the invariant part of the Lagrangian whereas the symmetry breaking part is treated as a rather weak perturbation. A dynamical mechanism of the nucleon mass generation through the mechanism of spontaneous breaking of the chiral symmetry was elaborated in independent works by Valentin Vaks and Anatoly Larkin [273] and Yoichiro Nambu [218], applying superconductivity methods to a Heisenberg type four-fermion two-component Lagrangian.
We should certainly mention a role played by the spontaneous symmetry breaking in the unified electroweak theory of Abdul Salam, Sidney Glashow and Steven Weinberg, and the role played by the Higgs scalar boson in it [102, 101, 240, 241, 297]. The study of the properties of superdense matter described by unified gauge theories began in 1972 with the work of David Kirzhnits [146], who showed that the classical scalar field responsible for symmetry breaking disappears at a high enough temperature. In 1974 first grand unified theory was proposed by Howard Georgi and Sidney Glashow [98].
Now let us come to the subject of this review, i.e., discussion of yet considered as exotica, possibilities of existence of various abnormal nuclear systems. Soon after prediction of quarks by Murrey Gell-Mann and George Zweig in 1964, Dmitry Ivanenko and Dmitry Kurdgelaidze suggested that a neutron star may transform in a quark star [128], see also [127]. In 1971 Arnold Bodmer proposed that also atomic nuclei may undergo a transition in a superdense quark state [34].
In 1971 Arkady Migdal put forward the idea that ground state of atomic nuclei may contain a pseudoscalar condensate ($0^{-}$) of pions. He demonstrated occurrence of the pion condensation in a sufficiently deep potential well in scalar and electric fields forming deep potential wells for pions and suggested possibility of superdense stable or metastable nuclei glued by the pion condensate [204]. Three types of abnormal $\pi$ condensate nuclei: superdense, neutron and supercharged ones were discussed in [214, 211], more details see in reviews [215, 210].
In 1974 Tsung-Dao Lee and Gian-Carlo Wick suggested possible existence of a superdense state of nuclear matter characterized by a value of the scalar (e.g., $\sigma$) field different from its mean value in atomic nuclei [181, 179]. This idea was reviewed in [182, 183]. The simplest variant of the model contains only two fields, the neutral scalar meson field and the isospinor nucleon field. As the $\pi$ condensate nuclei, the $\sigma$ abnormal nuclear state can be either stable or metastable. In the former case the normal nuclei after a while would undergo the phase transition to the new state. In first works the mean field approximation was used in spite of the assumed strong interaction of the scalar meson with the nucleons. To somehow justify this assumption the authors noted analogy between the abnormal state and the Bose–Einstein condensation one. Then Lee and Margulies argued [176, 180] that quantum fluctuations may not change the statement about the existence of a super-dense state of nuclear matter.
At the workshop at Bear Mountain, to north from New York City, on “BeV/nucleon collisions of heavy ions” Lee presented his results on the possible vacuum reconfiguration [181] and noted [175] that “in order to produce the abnormal nuclear state, we must consider reactions in which (i) a large number of nucleons is involved and (ii) the nucleon density can be increased by a significant factor. One is, therefore, led to considerations of high energy collisions between heavy ions, say ${\rm U}+{\rm U}\to Ab+\dots$, where $Ab$ denotes the abnormal nuclear state.” The event at Bear Mountain was pivotal in the conception of heavy ion physics. 1974 there were the Synchrophasotron at the Joint Institute for Nuclear Research in Dubna and a high-energy heavy-ion accelerator, the Bevalac, being completed at the Lawrence Berkeley Laboratory in Berkeley. Ideas about possibility of the production of abnormal nuclei in heavy-ion collisions gave a push to active experimental investigations. One suggested also the look for collapsed quark nuclei [34] and superdense nuclei with pion condensate in heavy-ion collisions, see Refs. [7, 94]. The possibility of achieving a high degree of nuclear matter compression in shock waves, which could contribute to the formation of anomalously dense droplets in heavy ion collisions, was investigated in Refs. [250, 263]. Limits on the yields of abnormal nuclei in heavy-ion collisions and on their possible terrestrial concentrations were discussed in Refs. [124, 141, 14, 13, 13, 15]. The interpretation of the data collected from heavy-ion collisions so far does not require the presence of abnormal nuclei. Feasibly, one of reasons is that compressed nuclear matter is produced in heavy ion collisions at a higher temperature than the binding energy of abnormal nuclei. Nevertheless, further dedicated experimental and theoretical analysis is needed.
Cosmos was another domain to look for an anomalous state of matter. Yakov Zeldovich liked to claim that the Universe is a poor man’s accelerator: experiments don’t need to be funded, and all we have to do is to collect the experimental data and interpret them properly. Years later ideas of spontaneous symmetry breaking penetrated in cosmological models, cf. [186].
Probably in 1975, Migdal returned from a conference and told one of us (D.N.V.) that during his talk, in which he mentioned a possibility of the existence of anomalous supercharged nuclei with a pion condensate, Gell-Mann asked him about a role of electrons in such highly-charged systems. Migdal suggested to think about this problem. In works [212, 209] the authors developed a relativistic semiclassical treatment of the Dirac equation in deep electric potentials and the many-particle approximation for electrons occupying levels of the lower continuum. The occupation of the vacuum levels by electrons and negative muons in the deep electric potential well was named “a fermion condensation”. The presence of the upper bound on the electric potential was discussed by Berndt Müller and Johann Rafelski in [217]. Preparing this manuscript, we realized that Lee also discussed a role of electrons in abnormal nuclei. He wrote in Ref. [183] that “…the abnormal state can create $e^{+}e^{-}$ pairs. The $e^{+}$ will be sent to infinity, but a fair fraction of the $e^{-}$ will be kept within the abnormal nucleus. As $Z$ increases, the number of $e^{-}$ also increases. The interplay between the added Fermi energy of $e^{-}$ and the Coulomb energy may eventually bring the abnormal state to the point of instability when $A$ reaches $\approx 10^{4}$.” References [280, 275, 282] described the charged condensate pions and vacuum electrons in strong electric fields and suggested a possibility of existence of “nuclei-stars” of arbitrary size (up to the compact star size) glued by the charged $\pi$ condensate produced in the dense nuclear matter and the electrons produced from the vacuum.
In 1979 Mikhail Troitsky and Nikolai Chekunaev discussed in [269] that at densities above the nuclear saturation density, $n_{0}=0.16$ fm-3, the gain in the energy due to the pion condensation can exceed the energy loss associated with the transformation of nucleons into heavier baryon species like hyperons and $\Delta$ isobars. The similar idea was applied to quarks by Edward Witten in 1984. In Ref. [298], he argued that in deconfined quark matter, as the baryon density increases, it becomes energetically favorable to occupy the $s$-quark Fermi sea alongside the $u$- and $d$-quark Fermi seas. Moreover, he claimed that the proposed ‘strange matter’, which contains comparable numbers of $u$, $d$, and $s$ quarks may be absolutely stable, i.e. more stable than an iron atomic nucleus. The Migdal hypothesis of abnormal nuclei with a pion condensate, the Lee-Wick hypothesis of abnormal nuclei with a $\sigma$ condensate and the Witten strange matter hypothesis may play an important role in cosmology, compact stars, cosmic ray physics, and relativistic heavy-ion collisions. Developing Witten’s idea, Edward Farhi and Robert Jaffe examined the strange matter clusters with an intermediate baryon number $10^{2}<A<10^{7}$, which are in equilibrium with respect to weak interactions [80]. Later Alvaro De Rujula and Sidney Glashow suggested the possibility of the existence of quark strangelets, large-size nuclearites and strange stars of a larger mass (when gravity comes into play), see Refs. [64, 65]. The charge distribution in the (strangelets and nuclearites) and the strange stars was considered in [6] in a similar manner as was done for supercharged nuclei in Refs. [209, 280, 275]. The surface layer of strange objects (strangelets, quark nuclearites and strange stars) has typical thickness $\sim 10$ fm. Also it was suggested that there may exist stars with a hadronic shell of a macroscopic size and quark core –hybrid stars [260, 28].
@@ I Introduction
The axion quark nuggets (AQN) model was proposed in Refs. [302, 303, 301]. The latter model incorporates axion domain walls formed during the QCD phase transition. The domain walls act as stabilizing structures, alleviating the need for a first-order phase transition as in the original Witten model. A generic feature of the quark nuggets is the electrosphere. In case of antimatter axion quark nuggets the electrosphere of the $\sim 10^{-8}$ cm depth contains positrons. The axion quark nuggets behave as conventional dark matter components in the low density environment and become strongly interacting macroscopically large objects in the relatively high-density environment. The available energy due to matter-antimatter annihilation can be as high as $\sim 2$ GeV per baryon charge.
@@
Although the initial studies [181, 179, 214] discussed the possibility of the sigma or pion condensation occurring at densities both greater than and less than $n_{0}$, it soon became evident that the latter possibility was unrealistic. On the other hand various models of equation of state of the dilute nearly isospin-symmetric nuclear matter demonstrate the presence of region of the first-order phase transition and spinodal instability at nucleon densities $0.3n_{0}\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}n\stackrel{{\scriptstyle\scriptstyle<}}{{\phantom{}{}_{\sim}}}0.7n_{0}$ [235, 60, 27, 254, 39]. In this density region the compressibility of nuclear matter is negative and the scalar Landau-Migdal parameter of the Fermi liquid $f_{0}(n)$ is smaller than 1-1, that causes the Pomeranchuck instability. In [158] it was shown that the Pomeranchuck instability may lead to a condensation of an effective scalar field in a Fermi liquid. The presence of such condensate would remove instabilities in first and zeroth sounds. Owing to the condensate, a novel metastable state might exist at nuclear subsaturation density for $N\sim Z$, where $N$ is the number of neutrons and $Z$ is the number of protons. In [198] it was realized that the stability region remains also in case of significantly isospin-asymmetric matter.
@@
It was found in Refs, [289, 291] that a rapidly rotating drop of nuclear matter, being not bound in the rest frame, may become bound in the rotation frame because of the formation of a charged pion condensate giant vortex. This observation invites looking for the nuclear drops with high angular momenta.
A few words about the notion “abnormal” that we use in the title are in order. Often, a new concept seems ‘abnormal’ but is considered as normal after a while. Therefore, we should clarify beforehand what we mean by the term abnormal nuclear systems. Various definitions were used in the literature: the scalar-condensate abnormal nuclei; the pion-condensate abnormal nuclei and nuclei-stars; nuggets; strangelets; nuclearites; strange stars and hybrid stars; quark stars and some others. Different authors have given different meanings to these terms. We will basically follow the definitions introduced by pioneers in this field: The term abnormal nuclei was used by Migdal to describe hypothetical superdense (with a density larger than several $n_{0}$) stable or metastable nuclei glued together by a pion condensate as well as by Lee who applied it to hypothetical superdense nuclei glued together by a scalar condensate. Farhi and Jaffe in 1984 introduced term strangelets to describe hypothetical stable quark systems of a microscopic size bound more strongly than 57Fe, which should contain a similar number of up, down, and strange quarks. Hypothetical quark objects of a larger (macroscopic) size De Ruhula and Glashow in 1984 named nuclearites. Below all abnormal stable hypothetical nuclear systems of such a size will be called nuclearites. For such objects the gravity contributes still much weaker than nuclear forces. When the baryon number becomes $A\stackrel{{\scriptstyle\scriptstyle>}}{{\phantom{}{}_{\sim}}}10^{56}$, the gravity starts to contribute significantly to the binding, and we are dealing with nuclei-stars. For nuclei-stars consisting of the strange quark matter (bound even at smaller $A$) we will use the term strange stars. Compact stars with quark interiors (which are not necessarily self-bound without gravitational attraction) and hadron exteriors of a macroscopic size are usually called hybrid stars.
The existence of stable or metastable nuclear systems of various sizes would have numerous important consequences in physics. Therefore, it would be worthwhile to continue theoretical studies and, especially, experimental searches of such objects.
The paper has the following structure. In Sect. II we focus on the Lee-Wick model of abnormal nuclei with a scalar (e.g. a $\sigma$ field) condensate. We discuss the relationship between the Lee-Wick model and the standard RMF models, as well as the RMF models with $\sigma$-scaled hadron masses and coupling constants. Section III considers various aspects of the model of supercharged nuclei held together by light charged bosons, if the latter existed. In reality the role of the light boson could be played by an ordinary pion, which can become effectively lighter in a dense nuclear medium and/or in external fields, for example in rapidly rotating nuclear matter. In Sect. IV we discuss density dependence of the p-wave pion-nucleon-nucleon and pion-nucleon-$\Delta$ isobar interactions which may result in charged and neutral pion condensates and a possibility of existence of superdense abnormal nuclei as proposed by Migdal. These phenomena can be modeled using the $\sigma$ model. Then we turn our attention to the subtleties associated with the description of the s-wave pion-nucleon interaction and the s-wave pion condensation. In Sect. V we consider $\Delta$-resonance matter. The uncertainty in the $\Delta$ optical potential allows for a wide variation in its value. For certain values, metastable or even stable resonance matter and abnormal nuclei may exist. The concept of $\Delta$-resonance matter is somewhat similar to Witten’s idea of strange quark matter. In the latter case the occupation of the strange quark Fermi sea results in a decrease of the energy. The energy of the state also decreases with a decrease of the strange quark mass and the bag constant, the values of which are not well known. In Sect. VI we focus on the hypothesis of strangelets and strange and hybrid stars. Ordinary neutron stars and stars with exotic matter have different mass-radius relations, and their cooling histories are different. These topics will be discussed. Section VII presents the idea of the scalar field condensation and possible existence of dilute metastable nuclear states. A scalar condensate field, determined by the value of the Landau-Migdal parameter in the scalar channel, may appear owing to the Pomeranchuk instability occurring in a dilute nuclear matter. Thereby we apply the idea of Lee and Wick of the possible existence of dense abnormal nuclei held together by a scalar field. Then, we consider the possibility of the formation of a self-bound dilute Bose-Einstein condensate state, e.g.,of $\alpha$ particles in the $\beta$-equilibrated system described with the help of a scalar condensate field. Next, we discuss anomalous states in nonequilibrium nuclear systems. Section VIII discusses anomalies that defy conventional physics interpretations, part of which could be related to abnormal nuclear systems. Conclusions are drawn in Sect. IX.
@@
II The Lee-Wick model
In this section we present the idea of Lee and Wick in a nutshell. We consider the Lagrangian density of nucleons interacting with a scalar field $\phi$: